# Synthia: Novel Concept Design with Affordance Composition

### v2

@@ 4 Experiments
4.1 Experimental Setup
Datasets.
To train T2I models with our approach, we construct a dataset from two types of resources (more details in Appendix D.1): (1) Existing Concept Images: For each existing concept in our ontology, we collect 60 images from external platforms including Google Images and iStock. To ensure that images are object-centric and aligned with the concept, we filter out low-quality images using CLIP model (Radford et al., 2021). (2) Generated Novel Concept Images: With our affordance sampling, we uniformly sample 600 affordance pairs among 235K possible pairs for fine-tuning. For the test dataset, we select 500 affordance pairs among the ones not used for fine-tuning.
Foundation Models
We adopt Kandinsky3.0 (Vladimir et al., 2024) as the T2I backbone model; it generates images based on a given text prompt, with an optional negative text prompt to refine outputs. During fine-tuning, we incorporate the desired affordances as positive inputs, while using the existing concepts from the ontology as negative constraints. During the inference, we provide only the text prompts with desired affordances, “a new design that has functions of {desired affordances}.”. Training details are provided in Appendix D.3
Figure 3: Results of the relative automatic evaluation. We compare the quality of concepts generated from our models and baselines with ones generated from our data generation pipeline (§3.1). Numbers indicate the percentage (%) of baseline model wins, ties, and DALLE model wins.
Baselines Methods
We compare our proposed method against three baseline methods, which are Stable Diffusion Esser et al. (2024), Kandinsky Arkhipkin et al. (2023), and ConceptLab Richardson et al. (2024). While Stable Diffusion and Kandinsky are general T2I models, ConceptLab optimizes generation over diffusion before creative concept design. For a fair comparison, we fine-tune ConceptLab using the same training data as our method. In contrast, our framework directly fine-tuned the diffusion model, integrating the hierarchical visual ontology to enforce the design of a single, coherent concept that achieves multiple affordances. Details on the baselines can be found in Appendix D.2.
@@ 4.2 Evaluation Metrics
To automatically evaluate the performance of our proposed method, we design four novel metrics to assess the quality of the generated data:
@@
For all four metrics, we use absolute scores ranging from 1 to 5, with higher scores indicating better quality. However, since the scores for these metrics may be influenced by subjective interpretation, we also include a relative evaluation. Specifically, we present each generated image with its corresponding DALL-E generated image, and ask the automatic evaluator to compare and determine which is superior or if they are equally strong. This relative comparison ensures a more fair evaluation and reduces potential biases. We use GPT-4o (OpenAI, 2024) as the evaluation model to assess the metrics. The detailed evaluation prompts used are provided in Appendix B.2.
@@
To assess the quality of the generated concepts beyond automated evaluations, we conduct a human evaluation with 36 non-design expert annotators. Recruited from the university across diverse majors, they are provided with a detailed rubric using the same metrics and a 1-5 scale as the automated evaluations. We randomly sample 10 affordance pairs for four models, with each sample independently evaluated. This allows direct comparisons between human and automated scores, capturing nuanced aspects of evaluation quality. Details of the evaluation process are documented in Appendix C to ensure transparency.
Figure 4: Effectiveness of curriculum learning. Learning curves indicate that our curriculum learning approach enables the model to generate novel concepts with significantly higher absolute evaluation scores even in the early stage of training, compared to the random training. The X-axis represents training steps.
@@ 4.3 Results and Analysis
Automatic Evaluation
In Table 1, we compare Synthia against three existing T2I models using our evaluation metrics. For a fair comparison, we randomly sample 500 test pairs that are not used for training. From the results in Table 1, we observe that the Stable Diffusion model always maintains a high practicality but lower novelty. This aligns with our observation (Fig 1) that it tends to generate existing concepts rather than novel concepts. In addition, when it cannot generate an existing concept that satisfies all affordances in the text prompt, the Stable Diffusion model will generate multiple objects in an image without any fusion. This is also reflected in the low coherence scores for both concept and affordance levels. For the other baseline methods, Kandinsky-3 and ConceptLab show an increase in terms of novelty and coherence. However, they suffer from a reduction in practicality. Compared to all the models, our method Synthia achieves the best faithfulness, novelty, and coherence scores while maintaining high performance in practicality. These results reflect that finetuning with the curriculum strategy can successfully follow the text instruction, fuse various affordances, and generate novel concepts.
Human Evaluation
To assess the consistency of human evaluations, we computed inter-annotator agreement (IAA) between two independent raters, where ratings were considered in agreement if their absolute difference was $\leq 1$. IAA across all images was 67.5% with Cohen’s Kappa of 22.3%, where Novelty achieved the highest agreement at 70.9%, followed by Faithfulness (68.5%), Practicality (66.4%), and Coherence (64.1%). Additionally, the agreement between aggregated human ratings and automatic evaluations reached 91.25%, indicating that human annotators achieve a reasonable level of consistency while automatic evaluations closely align with human judgments. The human evaluation results consistently demonstrate that our model generates functionally coherent and visually novel concepts with outperforming scores on faithfulness, novelty, practicality, and coherence (Table 1). More results are included in Figure 8.
@@ 4.4 Ablation Studies
The Size of Fine-tuning Training Data
@@
In our experiment, We fine-tune the diffusion model using 600 affordance pairs as the training data. To investigate the impact of the training data size, we compare performance across different scales: training with 200, 400, 600, and 800 affordance pairs and using automatic evaluation. As shown in Figure 6, we find that the performance improves with larger datasets, and reaches the optimal point at 600. Across all four training sizes, our model consistently outperforms baseline methods.
Number of Positive Affordances
To evaluate the impact of the number of positive affordances in the input prompt, we also conduct experiments using 3 and 4 positive affordances and compare the performance of all methods. As shown in Table 3 and 4, while a slight performance drop occurs across all models as the number of affordances increases, our method consistently maintains high novelty and coherence, showing the generalization ability of Synthia to integrate multiple affordances.
Effectiveness of Affordance Sampling
@@
To examine the impact of affordance pair distance on novelty, we select $100$ pairs with the lowest and $100$ pairs with the highest distance scores from the test set. The automatic novelty scores for each group, shown in Table 5, demonstrate that all three three baseline methods achieve relatively low novelty scores for close affordance pairs, which indicates a tendency to generate existing concepts rather than novel designs. In contrast, our method always exhibits high novelty across various distances and outperforms the baseline models.
Effectiveness of Curriculum Learning
@@
In our framework, we incorporate a curriculum learning (CL) strategy by gradually increasing the difficulty of the training during the fine-tuning process. To examine the importance of this component, we also compare the performance with and without curriculum learning (Fig 4). Specifically, we train the diffusion model by randomly shuffling the training data and computing the absolute automatic evaluation results. As shown in Table 6, without curriculum learning, we observe a performance drop and results demonstrate the effectiveness of the CL.
5 Conclusions and Future Work
Text-to-image models have shown great potential in concept generation. In this work, we introduce a framework for novel concept design, which integrates concept ontology construction, data generation, and a T2I model contrastive training pipeline with curriculum learning technique. In addition, we propose a four-dimensional metric that evaluates the quality of generated concept images. Experimental results across three strong T2I models from both automatic and human evaluations demonstrate that our method significantly outperforms the competing baseline methods. Ablation studies also highlight the importance of our affordance sampling and curriculum learning techniques.
@@ Acknowledgement
This research is based upon work supported by U.S. DARPA ECOLE Program No. HR00112390060. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ B.1 Data Generation
B.2 Automatic Evaluation
@@
Prompt for Absolute Automatic Evaluation Please act as an impartial evaluator to assess the quality of a single concept image generated by an AI, based on the user’s requirements. Your evaluation should use the following three criteria, each scored on a scale of 1 to 5: Faithfulness: Evaluate how well the object aligns with the provided instructions, including its intended affordances and functionalities. Does the text and image together indicate that the object serves the purpose for which it was designed? Scoring: 5: Flawlessly combines all specified functionalities as per the instructions. Text and image work in harmony to demonstrate a well-designed and fully functional object. 4: Fulfills most instructions and intended functionalities, with only minor inconsistencies or missing details. The text and image are mostly aligned. 3: Partially fulfills the instructions. Some functionalities are present but not well-integrated or consistent. There may be a minor mismatch between text and image. 2: Struggles to meet the provided instructions, missing key functionalities or combining them poorly. Text and image may conflict. 1: Does not follow the instructions at all. Functionalities are completely missing, irrelevant, or nonsensical. Novelty: Assess the originality and innovation of the design. Does the object show an exciting, novel design that would surprise or intrigue users? Scoring: 5: Highly innovative, unique, and impressive. Inspires curiosity or excitement, making it highly desirable to explore. 4: Contains interesting and novel elements, showing clear creative thought and appeal. 3: Displays moderate novelty, with some unique features but remaining relatively conventional or uninspiring. 2: Shows limited novelty, with minimal creativity and overly simplistic or derivative design. 1: Entirely unoriginal and mundane, lacking any creativity and appearing common or uninspiring. Practicality: Evaluate the real-world applicability of the object. Does the design make sense for human use? Would it align with human preferences and be feasible for production? Scoring: 5: Extremely practical and human-centric. Highly functional, aligns perfectly with human preferences, and seamlessly fits into real-world scenarios. 4: Mostly practical and applicable, with minor limitations that could be addressed to improve usability. 3: Somewhat practical but with notable flaws or unrealistic elements that may limit usability in real-world scenarios. 2: Largely impractical, with significant flaws or inconsistencies that make it unlikely to be useful. 1: Entirely impractical and unusable, failing to align with human preferences or real-world feasibility.
Prompt for Absolute Automatic Evaluation (continued) Coherence: This metric evaluates whether the image generated by the AI model contains only one primary object as instructed, focusing on the object’s clarity and functionality without the presence of additional, unintended objects. Scoring: 5: The image perfectly showcases one distinct object that aligns with the described attributes. There are no extraneous objects or elements that distract from the main object. 4: The primary object is clear and mostly isolated, but there may be minor elements in the background or periphery that do not significantly detract from the main object. 3: The main object is present and identifiable, but there are other elements in the image that somewhat distract from its clarity and functionality. 2: The image contains multiple objects where the main object is not clearly dominant or distinguishable from other unnecessary elements. 1: The image primarily features multiple objects, making it difficult to identify the intended single object; the composition is cluttered or entirely irrelevant to the instruction. Provide a score for each criterion, followed by a concise explanation justifying your ratings. Your final response should strictly follow this format:

### v4

@@ 4 Experiments
4.1 Experimental Setup
Datasets.
To train T2I models with our approach, we construct a dataset from two types of resources (more details in the Appendix B.2): (1) Existing Concept Images: For each existing concept in our ontology, we collect 60 images from external platforms including Google Images and iStock. To ensure that images are object-centric and aligned with the concept, we filter out low quality images using the CLIP model (radford2021learningtransferablevisualmodels). (2) Generated Novel Concept Images: With our affordance sampling, we uniformly sample 600 affordance pairs among 235K possible pairs for fine-tuning. For the test dataset, we select 500 affordance pairs among the ones not used for fine-tuning.
Baselines Methods.
We compare our proposed method against three baseline methods, which are Stable Diffusion esser2024scalingrectifiedflowtransformers, Kandinsky arkhipkin2023kandinsky, and ConceptLab richardson2024conceptlab. While Stable Diffusion and Kandinsky are general T2I models, ConceptLab optimizes generation over diffusion before creative concept design. For a fair comparison, we fine-tune ConceptLab using the same training data as our method. In contrast, our framework directly fine-tuned the diffusion model, integrating the hierarchical visual ontology to enforce the design of a single, coherent concept that achieves multiple affordances. Details on the baselines can be found in the Appendix B.3.
@@ 4.1 Experimental Setup
Implementation Details.
We adopt Kandinsky3.0 (vladimir-etal-2024-kandinsky) as the T2I backbone model; it generates images based on a given text prompt, with an optional negative text prompt to refine outputs. During fine-tuning, we incorporate the desired affordances as positive inputs, while using the existing concepts from the ontology as negative constraints. During inference, we provide only text prompts with desired affordances, “a new design that has functions of {desired affordances}.”. Training details are provided in the Appendix B.4
@@ 4.2 Evaluation Metrics
We adopt the LLM-as-a-Judge (zheng2023judging; kim2023prometheus) for automatic evaluation, due to its strong alignment with human judgments. We especially utilize a proprietary model, GPT-4o (openai2024gpt4ocard), capable of processing both visual and textual inputs. Building on this, we design four novel metrics to automatically assess the quality of the generated data:
@@
For all four metrics, we use absolute scores ranging from 1 to 5, with higher scores indicating better quality. However, since the scores for these metrics may be influenced by subjective interpretation, we also include a relative evaluation. Specifically, we present each generated image with its corresponding DALL-E generated image, and ask the automatic evaluator to compare and determine which is superior or if they are equally strong. This relative comparison ensures a more fair evaluation and reduces potential biases (detailed in the Appendix B.5).
@@
To assess the quality of the generated concepts beyond automated evaluations, we conduct a human evaluation with 36 non-design expert annotators. Recruited from the university across diverse majors, they are provided with a detailed rubric using the same metrics and a 1-5 scale as automated evaluations. We randomly sample 10 affordance pairs for four models, with each sample independently evaluated. This allows direct comparisons between human and automated scores, capturing nuanced aspects of evaluation quality. Details of the evaluation process are documented in the Appendix B.6 to ensure transparency.
Figure 3: Results of the relative automatic evaluation. We compare the quality of concepts generated from our models and baselines with ones generated from our data generation pipeline (§3.1). Numbers indicate the percentage (%) of baseline model wins, ties, and DALL-E model wins.
@@ 4.3 Results and Analysis
Automatic Evaluation.
We compare Synthia against three existing T2I models using our evaluation metrics. To ensure a fair comparison, we randomly sample 500 test affordance pairs not seen during training. As shown in the absolute evaluation results (Table 1), Stable Diffusion consistently achieves high practicality but lower novelty. This aligns with our observation that it tends to generate existing concepts rather than novel ones (Fig. 1). When it fails to retrieve an existing concept that satisfies all the affordances in the prompt, it often generates multiple disjoint objects without meaningful fusion, leading to lower coherence scores for both concept and affordance levels. In contrast, Kandinsky-3 and ConceptLab exhibit improved novelty and coherence compared to Stable Diffusion, but at the cost of reduced practicality. Their outputs are often more imaginative but less grounded in affordances. Our method, Synthia, outperforms all baselines in faithfulness, novelty, and coherence, while also maintaining strong performance in practicality. These findings are reinforced by the relative evaluation results (Fig. 3 and 7), where Synthia consistently achieves higher win rates against DALL-E across all metrics compared to baselines. Notably, Synthia even surpasses DALL-E in overall preference, indicating that its generated concepts are perceived as more innovative and functionally coherent. These results demonstrates that fine-tuning with our curriculum strategy enables Synthia to better follow textual instructions, effectively integrate multiple affordances, and generate high-quality novel concepts.
Human Evaluation.
To assess the consistency of human evaluations, we compute inter-annotator agreement (IAA) between two independent raters, considering ratings to be in agreement if their absolute difference is $\leq 1$. Overall IAA across all images was 67.5%, with a Cohen’s Kappa of 22.3%. Among the evaluation criteria, Novelty shows the highest agreement (70.9%), followed by Faithfulness (68.5%), Practicality (66.4%), and Coherence (64.1%) (Fig. 8). Additionally, the agreement between aggregated human ratings and automatic evaluations reached 91.25%, suggesting strong alignment between human judgments and our automated evaluation framework. These results reinforce that human annotators demonstrate a reasonable level of consistency, and that our automatic evaluation metrics reliably reflect human preferences. Human evaluation further confirms that our model consistently generates functionally coherent and visually novel concepts, achieving superior performance across all metrics (Table 1).
@@ 4.4 Ablation Studies
The Size of Training Data.
Figure 4: Ablation with different number of training data. We show the absolute automatic evaluation results of Synthia trained with different number of data.
@@
In our experiment, we fine-tune the diffusion model using 600 affordance pairs as training data. To investigate the impact of training data size, we compare model performance across different dataset scales: 200, 400, 600, and 800 affordance pairs, evaluated using our automatic evaluation protocol. As shown in Fig. 4, performance improves as the dataset size increases, reaching its peak at 600 pairs. Notably, our model consistently outperforms all baseline methods across all training sizes, demonstrating the efficacy of our approach.
Effectiveness of Affordance Sampling.
@@
To examine how the distance between affordance pairs affects concept novelty, we sample $100$ test pairs with the closest and $100$ with the farthest distances. The automatic novelty scores for each group, shown in Table 6, demonstrate that all three baseline methods achieve relatively low novelty scores for closely related affordance pairs, indicating a tendency to replicate existing concepts rather than generating novel designs. In contrast, our method consistently exhibits high novelty across various distances, outperforming all baselines and demonstrating its ability to generate creative and diverse concepts regardless of affordance similarity.
Effectiveness of Curriculum Learning.
Figure 5: Effectiveness of curriculum learning. We show learning curves of Synthia with different training methods. The X-axis represents training steps.
@@
Our framework incorporates a curriculum learning (CL) strategy that gradually increases training difficulty during the fine-tuning process. To assess its effectiveness, we compare model performance with and without CL. Specifically, we train a base T2I model by randomly shuffling the training data and evaluate both settings using the absolute automatic evaluation protocol. As shown in the learning curve (Fig. 5), our CL approach leads to significantly higher scores even in the early stages of training compared to random training, highlighting its role in accelerating learning and guiding the model toward generating high-quality novel concepts. In contrast, training with CL results in a noticeable performance drop (Table 7), further demonstrating the importance of CL in our framework.
Number of Positive Affordances.
To evaluate the impact of increasing the number of positive affordances in the input prompt, we conduct additional experiments using prompts with three and four affordances. As shown in Tables 4 and 5, all methods experience a slight performance drop as input complexity increases. However, Synthia consistently maintains high scores in novelty and coherence, showing strong generalization to more complex affordance combinations—even though it is trained only on two affordance pairs. This result highlights the scalability of our method in generating high-quality, functionally coherent, and visually novel concepts without additional fine-tuning, showcasing its robustness and generalization capabilities. Moreover, Synthia shows strong performance using only 600 affordance pairs with curriculum learning, accelerating training and enhancing generalization to novel affordance configurations.
Model Faithfulness Novelty Practicality Coherence
Stable Diffusion $3.71$ $3.78$ $3.37$ $3.37$
Kandinsky3 $3.28$ $4.07$ $3.08$ $4.08$
ConceptLab $3.33$ $4.57$ $3.03$ $4.04$
DALL-E 4.08 $4.32$ $3.32$ $4.56$
Synthia (Ours) $4.01$ 4.57 3.41 4.78
Table 2: Evaluation Results with GPT-4o-mini model.
Evaluation Model Selection.
We further conducted experiments using different models, especially GPT-4o-mini, applying the same evaluation prompts and detailed criteria (Table 2). We observe consistent agreement in the results across different GPT-4o checkpoints, with our model consistently outperforming existing baselines across all metrics, even surpassing DALL-E model in novelty, practicality, and coherence, where DALL-E achieves 4.27, 3.21, and 4.52, respectively, when evaluated with GPT-4o-2024-08-06 (Faithfulness 4.02).
5 Conclusion
In this work, we take a step toward affordance-aware generative design by tackling a critical gap in existing text-to-image models—their inability to synthesize visually novel and functionally meaningful designs. With our model, Synthia, we move beyond stylistic variation to purposeful creation, incorporating a hierarchical concept ontology with curriculum-guided fine-tuning to teach models how to compose affordances into unified, novel concepts. Our proposed four-dimensional evaluation metrics capture how well generated concepts align with real-world design principles. Through both human and automatic evaluations, Synthia demonstrates clear superiority in designing imaginative but functionally grounded concepts. By bridging creativity with utility, this work advances AI-driven generative design.
@@ Acknowledgement
This research is based upon work supported by U.S. DARPA ECOLE Program No. #HR00112390060 and an Office of Naval Research (ONR) award #N00014-23-1-2780. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ B.1 Data Generation
To construct high-quality novel concept designs for each sampled affordance pair, we first generate candidate descriptions that satisfy the desired functionalities while ensuring visual novelty relative to existing concepts. Specifically, we prompt GPT-4o (openai2024gpt4ocard) to produce 10 diverse image captions for each affordance pair, describing novel concepts that fulfill the given affordances and differ in appearance from known objects. We then use these captions to generate 10 corresponding images using DALL-E (ramesh2021zeroshottexttoimagegeneration), resulting in a set of pseudo novel concepts.
@@
B.2 Dataset
  • Existing Concept Images: For each existing concept in our ontology, we collect a dataset of 60 images from external platforms including Google Images and iStock. To ensure that the dataset is object-centric and minimizes noise, we filter out low-quality images using CLIP model (radford2021learningtransferablevisualmodels). Specifically, we compute the similarity between the image embeddings and the text embeddings of the "a photo of {concept name}", selecting top-5 images with the highest similarity scores for each concept used as negative constraints.
  • Generated Novel Concept Images: With our affordance sampling, we uniformly sample 600 affordance pairs among 82K possible pairs for fine-tuning. For the test dataset, we select 500 affordance pairs among the ones not used for fine-tuning. We use the generated images from the sampled affordances for fine-tuning and evaluation. The overall statistics can be found in Table  3.
Dataset Fine-tuning Inference
Existing Generated Generated
# Concepts 772 600 500
# Images 3860 1800 500
Table 3: Statistics of the datasets.
B.3 Baseline Methods
  • Stable Diffusion (esser2024scalingrectifiedflowtransformers) is a strong baseline model for high-fidelity image synthesis, which is built on a diffusion-based framework. In this work, we leverage the pre-trained stable-diffusion-3.5-large model as the foundation model for the text-to-image task to generate a novel concept. Due to the limited context window length, we input only the positive constraints that contain the desired affordances, omitting the negative constraints associated with existing concepts.
  • Kandinsky (arkhipkin2023kandinsky; vladimir-etal-2024-kandinsky) model serves as another strong baseline model for comparison. We utilize the pre-trained Kandinsky 3.0 model without any finetuning as the baseline, aligning it with the foundation model used in our proposed framework. This approach ensures a consistent starting point for evaluation and a fair comparison. This baseline allows us to effectively demonstrate the impact of our proposed training framework by comparing performance before and after the fine-tuning process.
  • ConceptLab (richardson2024conceptlab) is a state-of-the-art framework designed for creative concept generation, which leverages an innovative approach that formulates the generation problem as an optimization process over the output space of the diffusion prior. It adopts a similar input format to the one used in our setting. We follow the finetuning process in the original framework, applying our training data to generate novel concepts. We then compare the generation quality during the inference time.
B.4 Hyperparameter Settings
To fine-tune the diffusion model with the constructed dataset from DALL-E, we use the optimizer AdamW (loshchilov2019decoupledweightdecayregularization) with a learning rate of $\eta=\{5*10^{-6},10^{-6},5*10^{-7}\}$ to avoid the catastrophic forgetting problem. During the curriculum learning, we split the dataset into three groups based on their difficulty and trained 20 epochs for each group. For the weight factor in the triplet loss, we set $\gamma=\{0,0.2,0.5,0.8,1\}$. We finetune the UNet part in the pre-trained model and freeze the weights of other components in the diffusion model. The training takes about 1 GPU hour with NVIDIA A100 GPU.
B.5 Automatic Evaluation
B.5.1 Absolute Automatic Evaluation
@@
Prompt for Absolute Automatic Evaluation Please act as an impartial evaluator to assess the quality of a single concept image generated by an AI, based on the user’s requirements. Your evaluation should use the following three criteria, each scored on a scale of 1 to 5: Faithfulness: Evaluate how well the object aligns with the provided instructions, including its intended affordances and functionalities. Does the text and image together indicate that the object serves the purpose for which it was designed? Scoring: 5: Flawlessly combines all specified functionalities as per the instructions. Text and image work in harmony to demonstrate a well-designed and fully functional object. 4: Fulfills most instructions and intended functionalities, with only minor inconsistencies or missing details. The text and image are mostly aligned. 3: Partially fulfills the instructions. Some functionalities are present but not well-integrated or consistent. There may be a minor mismatch between text and image. 2: Struggles to meet the provided instructions, missing key functionalities or combining them poorly. Text and image may conflict. 1: Does not follow the instructions at all. Functionalities are completely missing, irrelevant, or nonsensical. Novelty: Assess the originality and innovation of the design. Does the object show an exciting, novel design that would surprise or intrigue users? Scoring: 5: Highly innovative, unique, and impressive. Inspires curiosity or excitement, making it highly desirable to explore. 4: Contains interesting and novel elements, showing clear creative thought and appeal. 3: Displays moderate novelty, with some unique features but remaining relatively conventional or uninspiring. 2: Shows limited novelty, with minimal creativity and overly simplistic or derivative design. 1: Entirely unoriginal and mundane, lacking any creativity and appearing common or uninspiring. Practicality: Evaluate the real-world applicability of the object. Does the design make sense for human use? Would it align with human preferences and be feasible for production? Scoring: 5: Extremely practical and human-centric. Highly functional, aligns perfectly with human preferences, and seamlessly fits into real-world scenarios. 4: Mostly practical and applicable, with minor limitations that could be addressed to improve usability. 3: Somewhat practical but with notable flaws or unrealistic elements that may limit usability in real-world scenarios. 2: Largely impractical, with significant flaws or inconsistencies that make it unlikely to be useful. 1: Entirely impractical and unusable, failing to align with human preferences or real-world feasibility. Coherence: This metric evaluates whether the image generated by the AI model contains only one primary object as instructed, focusing on the object’s clarity and functionality without the presence of additional, unintended objects. Scoring: 5: The image perfectly showcases one distinct object that aligns with the described attributes. There are no extraneous objects or elements that distract from the main object. 4: The primary object is clear and mostly isolated, but there may be minor elements in the background or periphery that do not significantly detract from the main object. 3: The main object is present and identifiable, but there are other elements in the image that somewhat distract from its clarity and functionality. 2: The image contains multiple objects where the main object is not clearly dominant or distinguishable from other unnecessary elements. 1: The image primarily features multiple objects, making it difficult to identify the intended single object; the composition is cluttered or entirely irrelevant to the instruction. Provide a score for each criterion, followed by a concise explanation justifying your ratings. Your final response should strictly follow this format: